OpenStudy (anonymous):

Which of the quadratic functions has the narrowest graph!? 1.)y=1/8x^2 2.)y=2x^2 3.)y=-x^2 4.)y=1/6x^2

OpenStudy (anonymous):

@Zale101 not sure i'd say y=-x^2 or y=1/6x^2? what do you think!:)

OpenStudy (zale101):

y=1/8x^2 is the narrowest of all of them

OpenStudy (anonymous):

really why so? looks wider

OpenStudy (zale101):

I graphed it and it looks wider

OpenStudy (zale101):



OpenStudy (zale101):

y=1/6x^2 is the second widest